OpenStudy (anonymous):

triangle RST~triangle JKL RS=8cm JK=12cm perimeter of triange RST =32cm what is the perimeter of triangle JKL?

Directrix (directrix):

If two triangles are similar, then the ratio of corresponding sides (scale factor) is the same as the ratio of the perimeters of the two triangles. Sides RS and JK are corresponding sides. RS/JK = 8/12 = 2/3 2/3 = 32/p where p is the perimeter of Triangle JKL Solve: 2/3 = 32/p for p. @brindli4